OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (−3, −2) B. (−1, 3) C. (1, 5) D. (−4, 0)

OpenStudy (ayeedomo):

What system of equations?

OpenStudy (anonymous):

ill post it right now

OpenStudy (anonymous):



OpenStudy (anonymous):

hello

OpenStudy (ayeedomo):

Hang on I'm calculating XD

OpenStudy (anonymous):

thx

OpenStudy (ayeedomo):

I believe it's D: 2(-4)+(0)=-8 -8+0=-8 (-4)-(0)=-4 -4 - 0 = -4

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

do u want a medal

OpenStudy (ayeedomo):

Yeah It's the only pair that works with both equations. And sure :)

OpenStudy (anonymous):

can u help me wif more XD

OpenStudy (ayeedomo):

Sure lol

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3

OpenStudy (anonymous):



OpenStudy (ayeedomo):

C because: (-2) - 4(-2)=6 -2 - -8 = 6 -2 + 8 = 6 (two negative signs becomes positive) 3(-2) + (-2) =-8 -6 + -2 = -8